If at the moment t=0, the separation between the cars is 90m and the length of each car is 5m, the moment of time at which they completely cross each other is t=
 

Grade:12th pass

3 Answers

Arun
5 years ago
Dear student
 
Your question is incomplete. Please check and repost it. You can also attach an image. I will be happy to help you.
 
Regards
Arun (askIITians forum expert)
Shivam Agarwal
2 years ago
Vp = 10m/sec
V = 15m/sec
S = 90m
length = 5m
Relative speed = 10+15= 25m/sec
Distance = 2×length+S = 2×5+90= 100m
Time = Distance/relative speed = 100/25 = 4sec
Prana
one year ago
Total distance = (2l+90) 
L is given 5                     l is given as 5
Distance is 100 m on solving 
Velocitt of p w. R. T Q is 25 m/s on solving displacement upon time 
T distance/velocity I. E 100m/25m/s = 4 sec